# Mambular: A Sequential Model for Tabular Deep Learning

### v1

@@ Appendix B Results
Mambular 0.018 $\pm$ 0.000 0.452 $\pm$ 0.043 0.167 $\pm$ 0.011 0.628 $\pm$ 0.010 0.035 $\pm$ 0.005 0.132 $\pm$ 0.020 0.027 $\pm$ 0.002
Appendix C Default Model Hyperparameters
In the following, we describe the default model parameters used for all the neural models. We based our choices on the literature to ensure meaningful and high-performing parameters by default. Additionally, except for MambaTab, we were able to reproduce results (on average) from popular studies, such as Gorishniy et al., (2021) and Grinsztajn et al., (2022). While most larger benchmark studies perform extensive hyperparameter tuning for each dataset, analyzing these results (Grinsztajn et al.,, 2022; Gorishniy et al.,, 2021) shows that most models already perform well without tuning, as also found by McElfresh et al., (2024). Furthermore, the results suggest that performing hyperparameter tuning for all models does not change the ranking between the models, since most models benefit from tuning to a similar degree. Thus, we have collected informed hyperparameter defaults which we list in the following. The hyperparameters such as learning rate, patience and number of epochs are shared among all models for a more consistent approach.
@@ Appendix C Default Model Hyperparameters
Lastly, we follow the Github implementation from Ahamed and Cheng, 2024a https://github.com/basf/mamba-tabular where the sequence is flipped and the SSM iterates over the number of observations instead of the pseudo-sequence length of 1.
@@
Mambular 0.927 $\pm$ 0.006 0.928 $\pm$ 0.002 0.856 $\pm$ 0.004 0.795 $\pm$ 0.011 0.917 $\pm$ 0.003
MambaTab
As described in the main part, MambaTab (Ahamed and Cheng, 2024a, ) significantly differs from Mambular, since it is not a sequential model. To achieve the presented results from MambaTab, we have followed the provided implementation from the authors retrieved from https://github.com/Atik-Ahamed/MambaTab. It is worth noting, however, that MambaTab benchmarks the model on a lot of smaller datasets. 50% of the benchmarked datasets have not more than 1000 observations. Additionally, the provided implementation suggests, that MambaTab does indeed not iterate over a pseudo sequence length of 1, but rather over the number of observations, similar to a TabPFN (Hollmann et al.,, 2022). We have also tested this version, denoted as MambaTabT but did not find that it performs better than the described version. On the Adult dataset, our achieved result of 0.901 AUC on average is very similar to the default results reported in Ahamed and Cheng, 2024a with 0.906. The difference could be firstly due to us performing 5-fold cross validation and secondly different seeds in model initialization.
Models BA $\uparrow$ AD $\uparrow$ CH $\uparrow$ FI $\uparrow$ MA $\uparrow$
@@
Appendix D California Housing
Given the results for the California Housing dataset for mambular, we further analyzed the dataset. Below are more results for Mambular with random shuffling. Again we can see the the position of Latitude and Longitude significantly impact model performance. We generally observe, that when more than 3 variables are between these two variables, this negatively impacts performance. Having these two right after each other, independent of the order, does not significantly impact the performance. However, we do find that the best model is the one where these two variables are right at the end.
Table 25: Analysis of results for CA Housing
Model CA $\downarrow$ $p$-value Ordering
[…]
0.174 $\pm$ 0.009 0.304 [Po, Hh, MI, OP, LO, LA, MA, TR, TB]
0.195 $\pm$ 0.012 0.005 [LO, MA, TR, TB, Po, Hh, MI, OP, LA]
0.197 $\pm$ 0.010 0.002 [MA, LO, TR, TB, Po, Hh, MI, LA, OP]
0.188 $\pm$ 0.010 0.014 [MA, TR, LO, TB, Po, Hh, LA, MI, OP]
0.178 $\pm$ 0.010 0.137 [MA, LO, LA, TR, TB, Po, Hh, MI, OP]
0.177 $\pm$ 0.008 0.142 [MA, TR, TB, Po, LA, LO, Hh, MI, OP]
0.178 $\pm$ 0.009 0.123 [LA, LO, MA, TR, TB, Po, Hh, MI, OP]
0.172 $\pm$ 0.070 0.420 [Hh, TB, Po, MI, MA, OP, LA, LO, TR]
0.177 $\pm$ 0.010 0.171 [LO, Po, OP, LA, MI, MA, TR, Hh, TB]
0.190 $\pm$ 0.010 0.009 [Hh, TB, LO, MI, Po, OP, TR, MA, LA]
D.1 Distributional Regression
Distributional regression describes regression that goes beyond the mere prediction of mean responses. In detail, all distributional parameters, e.g., mean, variance, kurtosis, and skewness, are jointly modeled and thus allow for a comprehensive understanding of the response distribution. This approach allows for a more detailed analysis by taking into account the entire distribution shape. It helps to make more accurate predictions and gain deeper insights into the variability and asymmetry within the data. For a comprehensive overview see Kneib et al., (2023).
CRPS
[…]
See Gneiting and Raftery, (2007) for more details.

### v2

@@ Appendix E MambaTab
Models BA $\uparrow$ AD $\uparrow$ CH $\uparrow$ FI $\uparrow$ MA $\uparrow$
Appendix F Default Model Hyperparameters
In the following, we describe the default model parameters used for all the neural models. We based our choices on the literature to ensure meaningful and high-performing parameters by default. Additionally, we were able to reproduce results (on average) from popular studies, such as Gorishniy et al., (2021) and Grinsztajn et al., (2022). While most larger benchmark studies perform extensive hyperparameter tuning for each dataset, analyzing these results (Grinsztajn et al.,, 2022; Gorishniy et al.,, 2021) shows that most models already perform well without tuning, as also found by McElfresh et al., (2024). Furthermore, the results suggest that performing hyperparameter tuning for all models does not change the ranking between the models, since most models benefit from tuning to a similar degree. Thus, we have collected informed hyperparameter defaults which we list in the following. The hyperparameters such as learning rate, patience and number of epochs are shared among all models for a more consistent approach.
@@ Appendix F Default Model Hyperparameters
Lastly, we follow the Github implementation from Ahamed and Cheng, 2024a where the sequence is flipped and the SSM iterates over the number of observations instead of the pseudo-sequence length of 1.
@@
Appendix G Additional Results
Additional results, including a comparison to MambaTab and the presented variants from the supplemental material above are given below.
RF 0.019 $\pm$ 0.001 0.461 $\pm$ 0.052 0.183 $\pm$ 0.008 0.485 $\pm$ 0.007 0.028 $\pm$ 0.006 0.121 $\pm$ 0.018 0.025 $\pm$ 0.002
LightGBM 0.019 $\pm$ 0.001 0.459 $\pm$ 0.047 0.171 $\pm$ 0.007 0.542 $\pm$ 0.013 0.039 $\pm$ 0.007 0.112 $\pm$ 0.018 0.023 $\pm$ 0.003
CatBoost 0.019 $\pm$ 0.000 0.457 $\pm$ 0.007 0.169 $\pm$ 0.006 0.583 $\pm$ 0.006 0.045 $\pm$ 0.006 0.106 $\pm$ 0.015 0.022 $\pm$ 0.001
NODE 0.019 $\pm$ 0.000 0.431 $\pm$ 0.052 0.207 $\pm$ 0.001 0.613 $\pm$ 0.006 0.045 $\pm$ 0.007 0.124 $\pm$ 0.015 0.026 $\pm$ 0.001
LinReg 0.115 $\pm$ 0.002 0.483 $\pm$ 0.055 0.365 $\pm$ 0.021 0.711 $\pm$ 0.006 0.830 $\pm$ 0.047 0.302 $\pm$ 0.033 0.289 $\pm$ 0.004
Mambular 0.018 $\pm$ 0.000 0.452 $\pm$ 0.043 0.167 $\pm$ 0.011 0.628 $\pm$ 0.010 0.035 $\pm$ 0.005 0.132 $\pm$ 0.020 0.025 $\pm$ 0.002
@@ Appendix G Additional Results
Mambular 0.927 $\pm$ 0.006 0.928 $\pm$ 0.002 0.861 $\pm$ 0.008 0.796 $\pm$ 0.013 0.917 $\pm$ 0.003
Table 26: Combined Ranking of Models for Regression and Classification Tasks
Models Regression Rank Classification Rank Overall Rank
XGBoost 5.14 $\pm$ 4.02 5.4 $\pm$ 4.51 5.25 $\pm$ 4.03
RF 5.00 $\pm$ 2.94 7.4 $\pm$ 3.36 6.00 $\pm$ 3.22
LightGBM 4.57 $\pm$ 2.07 3.4 $\pm$ 3.36 4.08 $\pm$ 2.61
CatBoost 4.00 $\pm$ 2.45 2.2 $\pm$ 1.10 3.25 $\pm$ 2.14
FT-Transformer 3.57 $\pm$ 2.51 4.6 $\pm$ 1.52 4.00 $\pm$ 2.13
MLP 10.86 $\pm$ 1.57 8.6 $\pm$ 3.36 9.92 $\pm$ 2.61
TabTransformer 10.80 $\pm$ 1.64 8.5 $\pm$ 1.91 9.78 $\pm$ 2.05
ResNet 8.14 $\pm$ 2.91 7.6 $\pm$ 3.05 7.92 $\pm$ 2.84
NODE 5.71 $\pm$ 2.93 7.6 $\pm$ 1.82 6.50 $\pm$ 2.61
Regression 13.57 $\pm$ 0.53 13.8 $\pm$ 0.45 13.67 $\pm$ 0.49
MambaTab 9.29 $\pm$ 2.56 11.6 $\pm$ 1.14 10.25 $\pm$ 2.34
MambaTabT 11.57 $\pm$ 1.40 12.2 $\pm$ 0.84 11.83 $\pm$ 1.19
MambaTab 7.11 $\pm$ 2.79 7.4 $\pm$ 1.67 7.25 $\pm$ 2.30
Mambular 4.00 $\pm$ 3.06 3.0 $\pm$ 1.22 3.58 $\pm$ 2.43